OpenStudy (anonymous):

Compare the graphs of the inverse variations y=-0.2/x and y=-0.3/x. please provide at least 3 comparisons

OpenStudy (coconutjj):

switch x and y around then solve for y

OpenStudy (anonymous):

@CoconutJJ You're talking about inverse functions, the question is about inverse variations =p Here is a graph of both, you can see they are quite alike: http://www.wolframalpha.com/input/?i=graph+y+%3D+-0.2%2Fx%2C+y+%3D+-0.3%2Fx You can tell that the graph of \(y = \frac{0.3}{x}\) is slightly farther away from zero compared to \(y = \frac{0.2}{x}\). Besides that they are both have shape of hyperbola, they both has an asymptote at x=0 where they approach \(\infty\) when approaching 0 from the right and \(-\infty\) from the left, and they both have a horizontal asymptote at y=0

OpenStudy (anonymous):

so how will it be written? x will be 1,2.3 and y will be -0.20, -0.4, -0.60?

OpenStudy (anonymous):

since x increases and y decreases

OpenStudy (anonymous):

I made a mistake. I analyzed \(y=\frac{0.2}{x}\) and \(y=\frac{0.3}{x}\) while it is \(y = -\frac{0.2}{x}\) and \(y=-\frac{0.3}{x}\). So the signs are swapped. That means they both approach \(-\infty\) when approaching zero from the right (instead of \(\infty\)) and approach \(\infty\) when approaching zero from the left (instead of \(-\infty\)). You mean x will be 1, 2, 3...? Well in that case we can take several samples of \(y=\frac{0.2}{x}\) $$ \begin{array}{c|cl} x&&y \\ \hline 0.125 & -\frac{0.2}{0.125} &= -1.6\\ \hline 0.25 & -\frac{0.2}{0.25} &= -0.8 \\\hline 0.5 & -\frac{0.2}{0.5} &= -0.4 \\\hline 1 & -\frac{0.2}{1} &= -0.2 \\\hline 2 & -\frac{0.2}{2} &= -0.1 \\\hline 3 & -\frac{0.2}{3} &= -0.0666 \dots \\\hline 4 & -\frac{0.2}{4} &= -0.05 \end{array} $$You can add more values for \(x\) (perhaps a couple of negative ones as well) or make a similar table for \(y=\frac{0.3}{x}\) Hope it helps =)

OpenStudy (anonymous):

thank you so so much!

OpenStudy (anonymous):

You're welcome =)